OpenStudy (anonymous):

Where does the Pythagorean Identity sin^2 θ+cos^2 θ=1 come from? And how does it relate to right triangles?

OpenStudy (anonymous):

@SithsAndGiggles @e.mccormick @texaschic101 @Luigi0210 @zepdrix @nincompoop Help please :)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[sin^2 \alpha =\dfrac{b^2}{c^2}\\cos^2\alpha = \dfrac{a^2}{c^2}\] add them you have \(sin^2 +cos^2 = \dfrac{b^2+a^2}{c^2}=\dfrac{c^2}{c^2}=1\)

OpenStudy (anonymous):

@troyandthor do you accept that \(a^2+b^2=c^2\) ?